# GS4: Generalizable Sparse Splatting Semantic SLAM

### v2

@@ Appendix B Reconstruction Results on ScanNet
| GS4 | 8.56 | 3.88 | 87.34 | 79.90 | 295k |
Appendix C Qualitative Comparison for Semantic Segmentation.
As illustrated in Fig. 5, our approach achieves superior semantic segmentation accuracy compared to the SGS-SLAM baseline. For example, in the first column of Fig.5, our semantic rendering provides a more accurate representation of the desks, chairs, and night tables than SGS-SLAM.
@@
Appendix D Zero-Shot Results on NYUv2
@@
| PSNR | 20.86 | 20.59 | 22.82 | 22.09 | |
| SSIM | 0.870 | 0.862 | 0.829 | 0.854 | |
| GS4 (Ours) | LPIPS | 0.230 | 0.251 | 0.322 | 0.268 |
| GS Num | 257k | 206k | 371k | 278k |
Appendix E Zero-Shot Results on TUM RGB-D

### v3

@@ Appendix F Reconstruction Results on ScanNet
| GS3LAM | 11.83 | 4.62 | 82.82 | 71.92 | 2154k |
| GS4 | 8.48 | 3.87 | 89.09 | 77.44 | 224k |
Appendix G Qualitative Comparison for Semantic Segmentation.
As illustrated in Fig. 5, our approach achieves superior semantic segmentation accuracy compared to the SGS-SLAM and GS3LAM baselines. For example, in the first column of Fig.5, our semantic rendering provides a more accurate representation of the desks, chairs, and night tables than baselines.
@@
Appendix H Zero-Shot Results on NYUv2
@@
| PSNR | 21.15 | 22.35 | 23.21 | 22.24 | |
| SSIM | 0.885 | 0.867 | 0.846 | 0.866 | |
| GS4 (Ours, 2nd stg) | LPIPS | 0.217 | 0.238 | 0.307 | 0.254 |
| GS Num | 273k | 228k | 392k | 298k |
Appendix I Zero-Shot Results on TUM RGB-D